OpenStudy (egracer):

A binomial has been partially factored, as shown below: x2 − 49 = (x + 7)(_____) Which of the following binomials represents the missing factor? (x + 42) (x − 42) (x + 7) (x − 7)

OpenStudy (campbell_st):

its the difference of 2 squares \[a^2 - b^2 = (a + b)(a - b)\]

OpenStudy (egracer):

so what would I do the figure this out... I'm a little confused on binomials

jimthompson5910 (jim_thompson5910):

look carefully at the formula campbell_st wrote out

jimthompson5910 (jim_thompson5910):

it might help to rewrite 49 as 7^2

OpenStudy (campbell_st):

well if you had \[x^2 - a^2 = (x -a)(x + ?)\]

OpenStudy (egracer):

x-7?

jimthompson5910 (jim_thompson5910):

correct x^2 - 49 = (x+7)(x-7)

OpenStudy (egracer):

Thank you!!

jimthompson5910 (jim_thompson5910):

np